Solve the initial value problem

(1+e^x)y' = e^(x-y) y(0) =1


1
Expert's answer
2020-07-23T18:04:05-0400

(ey)'=ex/(ex+1).

ey=ln(ex+1)+c.

if y(0)=1 then c=e-ln(2).

y=ln(ln(ex+1)+e-ln(2)).